# Comparing Specialised Small and General Large Language Models on Text Classification: 100 Labelled Samples to Achieve Break-Even Performance

### v3

@@ Acknowledgements
This work was partially supported by the projects funded by the European Union under the Horizon Europe: DisAI, GA No. 101079164, AI-CODE, GA No. 101135437; and by Modermed, a project funded by the Slovak Research and Development Agency, GA No. APVV-22-0414.
@@
Due to the significant computation costs incurred by the inference of larger language models (namely LLaMA-2, LLaMA-3, Mistral-7B and Zephyr-7B), instruction-tuning of the medium sized models (Mistral/Zephyr) and the additional other costs of the GPT model, we evaluate the models in a limited setting. The evaluation is done on a randomly selected set of 1 000 samples (following the practice in prior works Gao et al. (2021); Chang and Jia (2023); Sclar et al. (2024); Li and Qiu (2023); Köksal et al. (2023)). As the datasets are smaller (e.g., MRPC, BoolQ, CoLA, SNIPS, TREC) in many cases, this decision may not be as problematic, as the used 60:20:20 split leads to approximately the 1000 test samples. However, on the larger datasets (SST2, AG News and DB Pedia) this decision may skew the results. In addition, we run the GPT prompting and in-context learning only 10 times instead of 20 due to its costs.
@@ Appendix A Experimental Setup: Additional Details
Each experiment is repeated 100 times for fine-tuning and 20 times for the remaining approaches (due to the significant costs of inference or training of the larger language models) in order to observe and reduce the performance variance (as recommended by Gundersen et al. Gundersen et al. (2023) and Pecher et al. Pecher et al. (2024b)), since particularly in-context learning and fine-tuning with few samples were identified to produce results with significant variance due to effects of randomness Lu et al. (2022); Zhao et al. (2021); Zhang et al. (2022); Mosbach et al. (2021); Dodge et al. (2020). Each repeat uses the same fixed random seed for all models to guarantee deterministic and replicable results. The random seed covers all sources of randomness Gundersen et al. (2022b) – the split of data, selection of labelled samples, initialisation of models, order of data in training, non-deterministic operations (e.g., dropout), and selection of samples for in-context learning. At each point, we report the mean F1 macro and standard deviation.
@@ A.2 Approaches and Models Setup
For fine-tuning, we use the BERT (Devlin et al., 2019b) and RoBERTa (Liu et al., 2019) base models. We follow the typical setup and recommendations from related work (such as Mosbach et al. (2021); Dodge et al. (2020)), adding dropout layer with drop rate of 0.3 followed by a classification layer. We use learning rate of 1e-5 with AdamW optimiser with warmup and train the model until convergence using a maximum of 10 epochs, with variable batch size across different dataset subsets (starting at 4 and ending at 32 for the full dataset).
For prompting and in-context learning the Flan-T5 (Chung et al., 2024) base, GPT (4o-mini-2024-07-18 version) (Brown et al., 2020; Ouyang et al., 2022), LLaMA-2 (Touvron et al., 2023) 13B chat optimised model, LLaMA-3 (Dubey et al., 2024) 8B instruction optimised model, Mistral-7B (Jiang et al., 2023) and Zephyr-7B (Tunstall et al., 2023) models. For instruction-tuning we use the Flan-T5 (with full instruction-tuning), Mistral-7B and Zephyr-7B models (both instruction-tuned using LoRA (Hu et al., 2022)).
@@
For in-context learning, we use as many samples per class as the context-size of the models allow. As increasing this number after a certain point results in degraded performance, we always report the best performance achieved up until the dataset subset. As such, in some cases, we report the performance using 2-shot in-context learning classification even on full datasets, while in others, we report 100-shot. This is often the case with GPT model as it does not appear to benefit from increasing the number of in-context examples as much.
@@
All of the general models (Flan-T5, LLaMA-2, LLaMA-3, GPT, Mistral-7B and Zephyr-7B) are set to be deterministic (e.g., using no sampling, no beam search and using deterministic temperature) and set maximum number of tokens for generation to 10. In addition, due to the significant costs, we use 4-bit quantisation for the LLaMA-2 model. In case of Mistral-7B and Zephyr-7B, we use both the non-quantised version and the 4-bit quantised versions. For instruction-tuning we use only the 4-bit quantised versions of Mistral-7B and Zephyr-7B models due to the significant cost of training the models. The Flan-T5 model, the LLaMA-3 model and the GPT model are always used with full-precision.
Finally, for better results and comparison presentation, we group the individual models by size. We define the following 3 sizes: 1) small, which includes the Flan-T5 model; 2) medium, which includes the Mistral-7B and Zephyr-7B models; and 3) large, which includes the LLaMA-2, LLaMA-3 and GPT models.
@@ Appendix B Effect of Model Size on Comparison and Dataset Dependence
First of all, larger models do not consistently lead to better performance across different approaches. With the fine-tuning, we observe the expected behaviour. On lower number of labelled samples (up to $20-100$ depending on the dataset) the smaller BERT model is able to outperform the larger RoBERTa model. However, the larger model can achieve the highest performance on the different datasets using fewer labelled samples. As such, the individual break-even points appear earlier for the larger fine-tuning models. On the other hand, in case of prompting, in-context learning and instruction-tuning, the smaller models are often able to achieve better or similar performance than the larger counterparts even though the different in the number of parameters is large. For example, in-context learning with small models (Flan-T5) on the AG News and SST2 dataset outperforms in-context learning with medium (Mistral-7B and Zephyr-7B) or large (LLaMA and GPT) models, while on the BoolQ dataset the situation is opposite. Similarly, instruction-tuning with medium models (Mistral-7B and Zephyr-7B) models often leads to similar or lower performance than the instruction-tuning of the small (Flan-T5) model.

### v4

@@ Acknowledgements
This work was partially supported by the projects funded by the European Union under the Horizon Europe: DisAI, GA No. 101079164, AI-CODE, GA No. 101135437; by Modermed, a project funded by the Slovak Research and Development Agency, GA No. APVV-22-0414; and by the European Union NextGenerationEU through the Recovery and Resilience Plan for Slovakia under the project No. 09I03-03-V03-00020.
@@
Due to the significant computation costs incurred by the inference of larger language models (namely LLaMA-2, LLaMA-3, Mistral-7B and Zephyr-7B), instruction-tuning of the medium sized models (Mistral/Zephyr) and the additional other costs of the GPT-4 model, we evaluate the models in a limited setting. The evaluation is done on a randomly selected set of 1 000 samples (following the practice in prior works Gao et al. (2021); Chang and Jia (2023); Sclar et al. (2024); Li and Qiu (2023); Köksal et al. (2023)). As the datasets are smaller (e.g., MRPC, BoolQ, CoLA, SNIPS, TREC) in many cases, this decision may not be as problematic, as the used 60:20:20 split leads to approximately the 1000 test samples. However, on the larger datasets (SST2, AG News and DB Pedia) this decision may skew the results. In addition, we run the GPT-4 prompting and in-context learning only 10 times instead of 20 due to its costs.
@@ Appendix A Experimental Setup: Additional Details
Each experiment is repeated 100 times for fine-tuning and 20 times for the remaining approaches (due to the significant costs of inference or training of the larger language models) in order to observe and reduce the performance variance (as recommended by Gundersen et al. Gundersen et al. (2023) and Pecher et al. Pecher et al. (2024b)), since particularly in-context learning and fine-tuning with few samples were identified to produce results with significant variance due to effects of randomness Lu et al. (2022); Zhao et al. (2021); Zhang et al. (2022); Mosbach et al. (2021); Dodge et al. (2020). Each repeat uses the same fixed random seed for all models to guarantee deterministic and replicable results. The random seed covers all sources of randomness Gundersen et al. (2022a) – the split of data, selection of labelled samples, initialisation of models, order of data in training, non-deterministic operations (e.g., dropout), and selection of samples for in-context learning. At each point, we report the mean F1 macro and standard deviation.
@@ A.2 Approaches and Models Setup
For fine-tuning, we use the BERT (Devlin et al., 2019a) and RoBERTa (Liu et al., 2019) base models. We follow the typical setup and recommendations from related work (such as Mosbach et al. (2021); Dodge et al. (2020)), adding dropout layer with drop rate of 0.3 followed by a classification layer. We use learning rate of 1e-5 with AdamW optimiser with warmup and train the model until convergence using a maximum of 10 epochs, with variable batch size across different dataset subsets (starting at 4 and ending at 32 for the full dataset).
For prompting and in-context learning the Flan-T5 (Chung et al., 2024) base, GPT-4 (4o-mini-2024-07-18 version) (Brown et al., 2020; Ouyang et al., 2022), LLaMA-2 (Touvron et al., 2023) 13B chat optimised model, LLaMA-3 (Dubey et al., 2024) 8B instruction optimised model, Mistral-7B (Jiang et al., 2023) and Zephyr-7B (Tunstall et al., 2023) models. For instruction-tuning we use the Flan-T5 (with full instruction-tuning), Mistral-7B and Zephyr-7B models (both instruction-tuned using LoRA (Hu et al., 2022)).
@@
For in-context learning, we use as many samples per class as the context-size of the models allow. As increasing this number after a certain point results in degraded performance, we always report the best performance achieved up until the dataset subset. As such, in some cases, we report the performance using 2-shot in-context learning classification even on full datasets, while in others, we report 100-shot. This is often the case with GPT-4 model as it does not appear to benefit from increasing the number of in-context examples as much.
@@
All of the general models (Flan-T5, LLaMA-2, LLaMA-3, GPT-4, Mistral-7B and Zephyr-7B) are set to be deterministic (e.g., using no sampling, no beam search and using deterministic temperature) and set maximum number of tokens for generation to 10. In addition, due to the significant costs, we use 4-bit quantisation for the LLaMA-2 model. In case of Mistral-7B and Zephyr-7B, we use both the non-quantised version and the 4-bit quantised versions. For instruction-tuning we use only the 4-bit quantised versions of Mistral-7B and Zephyr-7B models due to the significant cost of training the models. The Flan-T5 model, the LLaMA-3 model and the GPT-4 model are always used with full-precision.
Finally, for better results and comparison presentation, we group the individual models by size. We define the following 3 sizes: 1) small, which includes the Flan-T5 model; 2) medium, which includes the Mistral-7B and Zephyr-7B models; and 3) large, which includes the LLaMA-2, LLaMA-3 and GPT-4 models.
@@ Appendix B Effect of Model Size on Comparison and Dataset Dependence
First of all, larger models do not consistently lead to better performance across different approaches. With the fine-tuning, we observe the expected behaviour. On lower number of labelled samples (up to 2010020-100 depending on the dataset) the smaller BERT model is able to outperform the larger RoBERTa model. However, the larger model can achieve the highest performance on the different datasets using fewer labelled samples. As such, the individual break-even points appear earlier for the larger fine-tuning models. On the other hand, in case of prompting, in-context learning and instruction-tuning, the smaller models are often able to achieve better or similar performance than the larger counterparts even though the different in the number of parameters is large. For example, in-context learning with small models (Flan-T5) on the AG News and SST2 dataset outperforms in-context learning with medium (Mistral-7B and Zephyr-7B) or large (LLaMA and GPT) models, while on the BoolQ dataset the situation is opposite. Similarly, instruction-tuning with medium models (Mistral-7B and Zephyr-7B) models often leads to similar or lower performance than the instruction-tuning of the small (Flan-T5) model.